OpenStudy (anonymous):

Solve the differential equation below, given the initial condition f(0) = 3. Express your answer using words or symbols. y=-5x^2

OpenStudy (lalaly):

wheres the DE

OpenStudy (anonymous):

the derivative is y'=-5x^2 my bad..

OpenStudy (lalaly):

\[\large{\frac{dy}{dx}=-5x^2}\]\[\large{dy=-5x^2dx}\]now integrate both sides\[\large{y=\frac{-5x^3}{3}+C}\]

OpenStudy (lalaly):

y(0)=3 \[\large{3=\frac{-5(0)}{3}+C}\] C=3 so\[y=\frac{-5x^3}{3}+3\]

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

f(x) = y f(0) is your x intercept